# A complete proof of the geometric Bombieri–Lang conjecture for ramified covers of abelian varieties

### v1

@@ 3 Boundedness of height on generic sequences
Let $X$ be as in the main Theorem 1.1. In this section we prove that the height of any generic sequence of rational points in $X(K)$ is bounded, which is the essential step in the proof of Theorem 1.1. The precise statement is as follows.
Proposition 3.1.
Let $K$ be the function field of a smooth projective curve $B$ over $\mathbb{C}$. Let $X$ be a projective variety of general type over $K$ with a finite morphism $f\colon X\to A$ to an abelian variety $A$ over $K$. Let $\{x_{n}\}_{n\geq 1}\subseteq X(K)$ be a generic sequence of rational points of $X$, i.e. any infinite subsequence of $\{x_{n}\}_{n\geq 1}$ is Zariski dense in $X$. Then the height of $\{x_{n}\}_{n\geq 1}$ is bounded.
The proof will be given in §3.4. In §3.1 we recall some Nevanlinna theory which will be used in the proof. In section §3.2 we prove a key lemma (Lemma 3.7) that is crucial to the proof. This was already mentioned in §1.2. In §3.3 we prove a simple result concerning positivity of a divisor on an abelian variety. The reader may directly read §3.4 first and come back to the previous three subsections when they need.
@@ 3.1 Preliminaries in Nevanlinna theory
Let $X$ be a projective variety over $\mathbb{C}$ and let $Z\subseteq X$ be a closed subscheme. Let $f\colon\mathbb{C}\to X$ be a holomorphic map such that $f(\mathbb{C})\nsubseteq Z$. We recall the definition of $N(r,f,Z)$ in [Yam1].
@@
Here ${\mathbb{D}}_{t}=\{z\in\mathbb{C}\big||z|<t\}$ is the open disk of radius $t$ with centre 0.
Moreover, when $f\colon{\mathbb{D}}_{R}\to X$ is a holomorphic map from an open disk ${\mathbb{D}}_{R}\ (R>0)$ to $X$ with $f({\mathbb{D}}_{R})\nsubseteq Z$, we can define the counting function $N(r,f,Z),N^{(k)}(r,f,Z)$ for any $r<R$ in the same way as above.
@@
    (Second Main Theorem, [Yam1, Theorem 3.1.1]) If $D$ is a reduced effective divisor on $A$, then
@@
    Here $\big|\big|$ means that the estimate holds for $r>0$ outside some exceptional set of finite Lebesgue measure. $T(r,f,D),T(r,f,L)$ are with respect to any smooth Hermitian metrics on ${\mathcal{O}}_{X}(D),L$ respectively. Similarly for (2) below.
@@
    ([Yam1, Corollary 2.5.8]) If $Z\subseteq A$ is a closed subscheme whose support has codimension greater than 1, then
@@
Actually, we will use Theorem 3.6 only for linear entire curves. In this case there is a simplified proof in [Yam3, pp. 255-257].
@@ 3.2 A key lemma
In this subsection we prove the following key lemma which will be used in §3.4 to prove the main result.
@@
$T_{(x,y,u,v)}(X_{(x,y)})\cong\{(u^{(1)},v^{(1)})\in{\mathbb{A}}^{2}\big|(x+2uv^{2})\cdot u^{(1)}+(y+2u^{2}v)\cdot v^{(1)}=0\}.$
@@
Note that we have fixed $(x_{n},y_{n})$. Therefore, as long as $v_{\infty}$ doesn’t satisfies the equation above, there must be $A\neq-1$ and $(u_{\infty}^{(1)},v_{\infty}^{(1)})\in T_{(0,0,0,v_{\infty})}(X_{s_{\infty},{\rm red}})$.
@@
Now we begin to prove Lemma 3.7. If $X_{s_{n}}$ is empty for all but finitely many $n\geq 1$, the lemma is trivial. Thus after taking a subsequence of $\{s_{n}\}_{n\geq 1}$, we assume $X_{s_{n}}$ is nonempty for all $n\geq 1$. Replacing $S$ by an irreducible component of Zariski closure of $\cup_{n\geq 1}\{s_{n}\}$ in $S$ if necessary, we may assume that the points $\{s_{n}\}_{n\geq 1}$ are Zariski dense in $S$. This implies that $\pi\colon X\to S$ is dominant. Replacing $X$ by $X_{\rm red}$, we may further assume that $X$ is reduced. Now by the generic reducedness (cf. [Sta, Lemma 054Z, Lemma 020I, Lemma 0578]), there exists a Zariski open dense subset $S_{0}$ of $S$ such that $X_{S_{0}}\to S_{0}$ has reduced fibres. Thus after taking a subsequence of $\{s_{n}\}_{n\geq 1}$, we may assume that $s_{n}\in S_{0}$ for all $n\geq 1$. In this case $X_{s_{n}}$ is reduced for all $n\geq 1$.
@@
By composing $p\big|_{Y}\colon Y\to X$ and $\pi\colon X\to S$, we get a (dominant) morphism $\pi\circ(p\big|_{Y})\colon Y\to S$. By the Raynaud–Gruson flattening theorem in [RG, Théorème 5.2.2], there exists a proper birational morphism $g\colon S^{\prime}\to S$ such that the strict transform $Y^{\prime}$ of $Y$ under $g$ is flat over $S^{\prime}$. Let $g_{Y}\colon Y^{\prime}\to Y$ be the natural morphism.
@@
Now for any point $y^{\prime}_{\infty}\in Y^{\prime}_{s^{\prime}_{\infty}}(\mathbb{C})$ and any Euclidean open neighbourhood $U_{y^{\prime}_{\infty}}$ of $y^{\prime}_{\infty}$ in $Y^{\prime}(\mathbb{C})$, by [BS, Theorem 2.12], the flatness of $Y^{\prime}\to S^{\prime}$ implies that the image of $U_{y^{\prime}_{\infty}}$ in $S^{\prime}(\mathbb{C})$ is open. In particular, it contains an (Euclidean) open neighbourhood of $s^{\prime}_{\infty}$ in $S^{\prime}(\mathbb{C})$. This means that $y^{\prime}_{\infty}$ is a limit point of the fibres $Y^{\prime}_{s^{\prime}_{n}}(\mathbb{C})(n\geq 1)$. In other words, there exists a sequence of points $y^{\prime}_{n}\in Y^{\prime}_{s^{\prime}_{n}}(\mathbb{C})$ tending to $y^{\prime}_{\infty}$ under Euclidean topology. Since $y^{\prime}_{\infty}$ is an arbitrary point on $Y^{\prime}_{s_{\infty}}(\mathbb{C})$, we have shown that every point of $Y^{\prime}_{s^{\prime}_{\infty}}(\mathbb{C})$ is a limit point of $Y^{\prime}_{s^{\prime}_{n}}(\mathbb{C})$. In short, we conclude that $Y^{\prime}_{s^{\prime}_{\infty}}$ is exactly the ‘limit’ of $\{Y^{\prime}_{s^{\prime}_{n}}\}_{n\geq 1}$.
@@
Step 3. Now we assume $\dim S=1$ and $X,S$ are normal. In particular $S$ is regular. We also assume that $X$ is irreducible (hence integral). Now since $\pi\colon X\to S$ is dominant, $\pi$ is flat. By [Sta, Lemma 0BRQ], there exists a finite surjective morphism $T\to S$ such that in the diagram
@@
Now for any sequence of vectors $\{(x_{n},v_{n})\in TX_{s_{n}}\}_{n\geq 1}$ tending to a tangent vector $(x_{\infty},v_{\infty})\in TX_{s_{\infty}}$, there is a unique sequence of vectors $\{(x_{n}^{\prime},v_{n}^{\prime})\in TX^{\prime}_{t_{n}}\}_{n\geq 1}$ corresponding to $\{(x_{n},v_{n})\in TX_{s_{n}}\}_{n\geq 1}$ under the natural identification $X^{\prime}_{t_{n}}\cong X_{s_{n}}(n\geq 1)$. We may assume that $v_{\infty}\neq 0$, otherwise the lemma is trivial. Then we can assume that $v_{n}\neq 0$ for all $n\geq 1$. Let $\overline{v}_{n},\overline{v}^{\prime}_{n}$ be the image of $v_{n},v^{\prime}_{n}$ in ${\mathbb{P}}T_{x^{\prime}_{n}}X^{\prime},{\mathbb{P}}T_{x_{n}}X$, respectively. Since ${\mathbb{P}}TX^{\prime}$ is compact, after taking a subsequence we can assume that $(x_{n}^{\prime},\overline{v}^{\prime}_{n})$ tending to an element $(x^{\prime}_{\infty},\overline{v}^{\prime}_{\infty})\in{\mathbb{P}}T_{t_{\infty}}(X^{\prime}_{t_{\infty}})$. This element maps to $(x_{\infty},\overline{v}_{\infty})$ under the tangent map. Choose some suitable real numbers $a_{n}$ such that $\{(x^{\prime}_{n},a_{n}v^{\prime}_{n})\}_{n\geq 1}$ has a subsequence tending to an element $(x^{\prime}_{\infty},v^{\prime}_{\infty})\in TX^{\prime}$ with $v^{\prime}_{\infty}\neq 0$. This implies that $(x_{n},a_{n}v_{n})$ has a subsequence tending to the image of $(x^{\prime}_{\infty},v^{\prime}_{\infty})$ in $TX$. Therefore $\overline{v}_{n}$ tends to the image of $\overline{v}^{\prime}_{\infty}$ in ${\mathbb{P}}T_{x_{\infty}}X$, which means that $v_{\infty}$ and the image of $v^{\prime}_{\infty}$ in $T_{x_{\infty}}X$ are proportional.
@@
3.3 A positivity lemma
@@
By [Se, Proposition 1.4], it suffices to show that for any curve $C\subseteq A$, $f_{*}(R)\cdot C>0$. We first assume $R$ is ample and effective. After a suitable translation, we may assume that $f(R)$ and $C$ intersect properly. Then $C$ does not intersect any irreducible component of $f(R)$ of codimension greater than 1. If $f(R)\cap C=\emptyset$, for any curve $C^{\prime}\subseteq f^{-1}(C)$, we have $R\cap C^{\prime}=\emptyset$, which contradicts to the ampleness of $R$. Therefore $C\cdot f_{*}(R)>0$. This proves the case when $R$ is ample and effective.
For general $R$, there exists a positive integer $m$ such that $mR$ is linearly equivalent to $R_{1}+R_{2}$, where $R_{1}$ is ample and effective and $R_{2}$ is effective. By the argument above, $f_{*}(R_{1})$ is ample on $A$. Moreover, $f_{*}(R_{2})$ is effective by definition, so it is also nef by the group structure of $A$ (cf. [Laz, Example 1.4.7]). Therefore $f_{*}(R_{1})+f_{*}(R_{2})$ is ample on $A$. By [Ful, Theorem 1.4], $f_{*}(R)$ is linearly equivalent to $f_{*}(R_{1})+f_{*}(R_{2})$, and thus it is also ample. ∎
@@
Since any effective divisor on an abelian variety is nef (cf. [Laz, Example 1.4.7]), we can easily prove that an effective divisor $D$ on an abelian variety is ample if and only if $D_{\rm red}$ is ample. Moreover, if $D_{1},D_{2}$ are two effective Cartier divisors on an abelian variety with ${\rm Supp}(D_{1})\subseteq{\rm Supp}(D_{2})$ and $D_{1}$ ample, then $D_{2}$ is also ample. This will be used in §3.4.
3.4 Proof of Proposition 3.1
@@
Now we prove Proposition 3.1 by contradiction. After taking normalization, we may assume that $X$ is normal. Assume to the contrary that the height of $\{x_{n}\}_{n\geq 1}$ is unbounded. After taking a subsequence, we can assume that the height of $\{x_{n}\}_{n\geq 1}$ tends to infinity as $n\to\infty$.
@@
First we repeat the following operations: Take $A_{1}=A,A_{0}={\rm Spec}\,K$ at the beginning. Then we have a finite map $f\colon X\to A_{1}\times A_{0}$. In each step, if there is a nontrivial quotient abelian variety $A^{\prime}$ of $A_{1}$ (i.e. a surjective homomorphism $A_{1}\to A^{\prime}$) such that the height of the image of $\{x_{n}\}_{n\geq 1}$ in $A^{\prime}$ under the composition $X\to A\to A^{\prime}$ does not tend to infinity, then after taking a subsequence of $\{x_{n}\}_{n\geq 1}$ we can assume that the image of $\{x_{n}\}_{n\geq 1}$ in $A^{\prime}$ has bounded height. Note that by our assumption in the previous paragraph, the height of $\{x_{n}\}_{n\geq 1}$ still tends to infinity. Change $A_{1},A_{0}$ to ${\rm ker}(A\to A^{\prime}),A^{\prime}\times A_{0}$ respectively. Composing $f$ with an isogeny $A\to A_{1}\times A_{0}$, we get a finite map $X\to A_{1}\times A_{0}$. This ends a step. Since the dimension of $A_{1}$ decreases strictly, the operations terminate after finitely many steps. Then we get a finite map $X\to A_{1}\times A_{0}$ such that the image of $\{x_{n}\}_{n\geq 1}$ in any nontrivial quotient abelian variety of $A_{1}$ has height tending to infinity, while the image of $\{x_{n}\}_{n\geq 1}$ in $A_{0}(K)$ has bounded height. Note that, by construction, the abelian variety $A_{1}$ is nontrivial.
@@
By Lang–Néron theorem (cf. [XY1, Theorem 3.4]), the image of $\{x_{n}\}_{n\geq 1}$ in $A_{0}(K)$ is contained in a finite union of translations of $A_{0}^{(K/\mathbb{C})}(\mathbb{C})$. Therefore the image of $X$ in $A_{0}$ is contained in a translation of $(A_{0}^{(K/\mathbb{C})})_{K}$. After composing $X\to A_{1}\times A_{0}$ with a translation, we may assume that the image of $X$ in $A_{0}$ is contained in $(A_{0}^{(K/\mathbb{C})})_{K}$. Let $T$ be the Zariski closure of the image of $\cup_{n\geq 1}\{x_{n}\}$ in $A_{0}^{(K/\mathbb{C})}$, which is a complex variety. Then the image of $X$ in $(A_{0}^{(K/\mathbb{C})})_{K}$ is exactly $T_{K}$. In conclusion, we get a finite morphism $X\to A_{1}\times_{K}T_{K}$, whose composition with the projection $A_{1}\times_{K}T_{K}\to T_{K}$ is surjective.
@@
From now on we focus on the finite morphism $X\to A_{1}\times_{K}T_{K}$. For simplicity, we replace the original $A$ by $A_{1}$ and replace the original $f$ by this map. In other words, we call this map $f\colon X\to A\times_{K}T_{K}$. After taking normalization for $T$, we may assume that $T$ is normal. Let $s_{n},t_{n}$ be the image of $x_{n}$ in $A(K),T_{K}(K)$ respectively, then $t_{n}\in T(\mathbb{C})$ is constant. Since $\{x_{n}\}_{n\geq 1}$ is generic in $X$ and $X\to T_{K}$ is surjective, $\{t_{n}\}_{n\geq 1}$ is also generic in $T$. Moreover, by our assumption, the image of $\{s_{n}\}_{n\geq 1}$ in any nontrivial quotient abelian variety of $A$ has height tending to infinity.
@@
We first consider the case when $f\colon X\to A\times_{K}T_{K}$ is not surjective. In this case, as mentioned in the introduction, we can still use the arguments in [XY2] combining with a flattening process.

### v2

@@ 3 Boundedness of height on generic sequences
The goal of this section is to prove Proposition 1.4, which asserts that the height of any generic sequence of rational points in $X(K)$ is bounded. Here $X$ can be a projective variety of general type which has a finite map to an abelian variety. This is the essential step in the proof of Theorem 1.1. Now we restate the proposition as follows.
Proposition 3.1 (Proposition 1.4).
In §3.1 we recall some Nevanlinna theory which will be used in the proof. As mentioned in §1.3, we will extend classical Nevanlinna theory to holomorphic maps on open discs and consider the limiting behaviour. In §3.2 we prove a key lemma (Lemma 3.7), which was already mentioned in §1.4 regarding the first difficulty.
The proof of Proposition 3.1 is carried out in the next three subsections. In §3.3 we perform an algebraic-geometric reduction, which resolves the latter two difficulties mentioned in §1.4. In §3.4 and §3.5 we realize the key counting argument in two stages: §3.4 obtains an upper bound for the numbers of tangency points between the branch divisor and the sections corresponding to the images in $A$ of certain rational points of $X$, or equivalently, a lower bound for their transversal intersection numbers. Finally in §3.5, we derive a contradiction via a reverse estimate.
We need to mention that the arguments in §3.3, §3.4 and §3.5 are coherent; although each subsection mainly focuses on its title, certain assumptions needed later are occasionally introduced early and will be explicitly noted. The reader may start directly from §3.3 and come back to the previous two subsections when they need.
@@ 3.1 Preliminaries in Nevanlinna theory
Let $X$ be a projective variety over $\mathbb{C}$ and let $Z\subseteq X$ be a closed subscheme. Let $f\colon\mathbb{C}\to X$ be a holomorphic map such that $f(\mathbb{C})\nsubseteq Z$. We recall the definition of $N(r,f,Z)$ in [YAM04].
@@
Here ${\mathbb{D}}_{t}=\{z\in\mathbb{C}\big||z|<t\}$ is the open disc of radius $t$ with centre 0.
Moreover, when $f\colon{\mathbb{D}}_{R}\to X$ is a holomorphic map from an open disc ${\mathbb{D}}_{R}\ (R>0)$ to $X$ with $f({\mathbb{D}}_{R})\nsubseteq Z$, we can define the counting function $N(r,f,Z),N^{(k)}(r,f,Z)$ for any $r<R$ in the same way as above.
@@
    (Second Main Theorem, [YAM04, Theorem 3.1.1]) If $D$ is a reduced effective divisor on $A$, then
@@
    Here $\big|\big|$ means that the estimate holds for $r>0$ outside some exceptional set of finite Lebesgue measure. $T(r,f,D),T(r,f,L)$ are with respect to any smooth Hermitian metrics on ${\mathcal{O}}_{A}(D),L$ respectively. Similarly for (2) below.
@@
    ([YAM04, Corollary 2.5.8]) If $Z\subseteq A$ is a closed subscheme whose support has codimension greater than 1, then
@@
Actually, we will use Theorem 3.6 only for linear entire curves. In this case there is a simplified proof in [YAM15b, pp. 255-257].
@@ 3.2 A key lemma
In this subsection we prove the following key lemma which will be used in §3.5.
@@
$T_{(x,y,u,v)}(X_{(x,y)})\simeq\{(u^{(1)},v^{(1)})\in{\mathbb{A}}^{2}\big|(x+2uv^{2})\cdot u^{(1)}+(y+2u^{2}v)\cdot v^{(1)}=0\}.$
@@
Note that we have fixed $(x_{n},y_{n})$. Therefore, as long as $v_{\infty}$ doesn’t satisfy the equation above, there must be $A\neq-1$ and $(u_{\infty}^{(1)},v_{\infty}^{(1)})\in T_{(0,0,0,v_{\infty})}(X_{s_{\infty},{\rm red}})$.
@@
Now we begin to prove Lemma 3.7. If $X_{s_{n}}$ is empty for all but finitely many $n\geq 1$, the lemma is trivial. Thus after taking a subsequence of $\{s_{n}\}_{n\geq 1}$, we assume $X_{s_{n}}$ is nonempty for all $n\geq 1$. Replacing $S$ by an irreducible component of Zariski closure of $\cup_{n\geq 1}\{s_{n}\}$ in $S$ if necessary, we may assume that the points $\{s_{n}\}_{n\geq 1}$ are Zariski dense in $S$. This implies that $\pi\colon X\to S$ is dominant. Replacing $X$ by $X_{\rm red}$, we may further assume that $X$ is reduced. Now by the generic reducedness (cf. [STA26, Lemma 054Z, Lemma 020I, Lemma 0578]), there exists a Zariski open dense subset $S_{0}$ of $S$ such that $X_{S_{0}}\to S_{0}$ has reduced fibres. Thus after taking a subsequence of $\{s_{n}\}_{n\geq 1}$, we may assume that $s_{n}\in S_{0}$ for all $n\geq 1$. In this case $X_{s_{n}}$ is reduced for all $n\geq 1$.
@@
By composing $p\big|_{Y}\colon Y\to X$ and $\pi\colon X\to S$, we get a (dominant) morphism $\pi\circ(p\big|_{Y})\colon Y\to S$. By the Raynaud–Gruson flattening theorem in [RG71, Théorème 5.2.2], there exists a proper birational morphism $g\colon S^{\prime}\to S$ such that the strict transform $Y^{\prime}$ of $Y$ under $g$ is flat over $S^{\prime}$. Let $g_{Y}\colon Y^{\prime}\to Y$ be the natural morphism.
@@
Now for any point $y^{\prime}_{\infty}\in Y^{\prime}_{s^{\prime}_{\infty}}(\mathbb{C})$ and any Euclidean open neighbourhood $U_{y^{\prime}_{\infty}}$ of $y^{\prime}_{\infty}$ in $Y^{\prime}(\mathbb{C})$, by [BS76, Theorem 2.12], the flatness of $Y^{\prime}\to S^{\prime}$ implies that the image of $U_{y^{\prime}_{\infty}}$ in $S^{\prime}(\mathbb{C})$ is open. In particular, it contains an (Euclidean) open neighbourhood of $s^{\prime}_{\infty}$ in $S^{\prime}(\mathbb{C})$. This means that $y^{\prime}_{\infty}$ is a limit point of the fibres $Y^{\prime}_{s^{\prime}_{n}}(\mathbb{C})(n\geq 1)$. In other words, there exists a sequence of points $y^{\prime}_{n}\in Y^{\prime}_{s^{\prime}_{n}}(\mathbb{C})$ tending to $y^{\prime}_{\infty}$ under Euclidean topology. Since $y^{\prime}_{\infty}$ is an arbitrary point on $Y^{\prime}_{s_{\infty}}(\mathbb{C})$, we have shown that every point of $Y^{\prime}_{s^{\prime}_{\infty}}(\mathbb{C})$ is a limit point of $Y^{\prime}_{s^{\prime}_{n}}(\mathbb{C})$. In short, we conclude that $Y^{\prime}_{s^{\prime}_{\infty}}$ is exactly the ‘limit’ of $\{Y^{\prime}_{s^{\prime}_{n}}\}_{n\geq 1}$.
@@
Step 3. Now we assume $\dim S=1$ and $X,S$ are normal. In particular $S$ is regular. We also assume that $X$ is irreducible (hence integral). Now since $\pi\colon X\to S$ is dominant, $\pi$ is flat. By [STA26, Lemma 0BRQ], there exists a finite surjective morphism $T\to S$ such that in the diagram
@@
Now for any sequence of vectors $\{(x_{n},v_{n})\in TX_{s_{n}}\}_{n\geq 1}$ tending to a tangent vector $(x_{\infty},v_{\infty})\in TX_{s_{\infty}}$, there is a unique sequence of vectors $\{(x_{n}^{\prime},v_{n}^{\prime})\in TX^{\prime}_{t_{n}}\}_{n\geq 1}$ corresponding to $\{(x_{n},v_{n})\in TX_{s_{n}}\}_{n\geq 1}$ under the natural identification $X^{\prime}_{t_{n}}\simeq X_{s_{n}}(n\geq 1)$. We may assume that $v_{\infty}\neq 0$, otherwise the lemma is trivial. Then we can assume that $v_{n}\neq 0$ for all $n\geq 1$. Let $\overline{v}_{n},\overline{v}^{\prime}_{n}$ be the image of $v_{n},v^{\prime}_{n}$ in ${\mathbb{P}}T_{x^{\prime}_{n}}X^{\prime},{\mathbb{P}}T_{x_{n}}X$, respectively. Since ${\mathbb{P}}TX^{\prime}$ is compact, after taking a subsequence we can assume that $(x_{n}^{\prime},\overline{v}^{\prime}_{n})$ tending to an element $(x^{\prime}_{\infty},\overline{v}^{\prime}_{\infty})\in{\mathbb{P}}T_{t_{\infty}}(X^{\prime}_{t_{\infty}})$. This element maps to $(x_{\infty},\overline{v}_{\infty})$ under the tangent map. Choose some suitable real numbers $a_{n}$ such that $\{(x^{\prime}_{n},a_{n}v^{\prime}_{n})\}_{n\geq 1}$ has a subsequence tending to an element $(x^{\prime}_{\infty},v^{\prime}_{\infty})\in TX^{\prime}$ with $v^{\prime}_{\infty}\neq 0$. This implies that $(x_{n},a_{n}v_{n})$ has a subsequence tending to the image of $(x^{\prime}_{\infty},v^{\prime}_{\infty})$ in $TX$. Therefore $\overline{v}_{n}$ tends to the image of $\overline{v}^{\prime}_{\infty}$ in ${\mathbb{P}}T_{x_{\infty}}X$, which means that $v_{\infty}$ and the image of $v^{\prime}_{\infty}$ in $T_{x_{\infty}}X$ are proportional.
@@
3.3 Step 1: algebraic-geometric reductions
Before proving Proposition 3.1, we first prove a simple lemma concerning the positivity of divisors on abelian varieties.
Recall that for a generically finite proper map $f\colon X\to Y$ between $K$-varieties, there is a push-forward homomorphism $f_{*}\colon{\rm Div}(X)\to{\rm Div}(Y)$ between the groups of Weil divisors. More precisely, for any prime Weil divisor $D$ of $X$, the image $D^{\prime}=f(D)$ is a closed subvariety of $Y$. If $\dim D^{\prime}<\dim D$, set $f_{*}(D)=0$. If $\dim D^{\prime}=\dim D$, then the function field $K(D)$ of $D$ is a finite extension of the function field $K(D^{\prime})$ of $D^{\prime}$. Set $f_{*}(D)=[K(D):K(D^{\prime})]\cdot D^{\prime}$. This extends linearly to a homomorphism $f_{*}\colon{\rm Div}(X)\to{\rm Div}(Y)$. By [FUL98, Theorem 1.4], $f_{*}$ descends to a homomorphism $f_{*}\colon{\rm Cl}(X)\to{\rm Cl}(Y)$ between Weil divisor class groups. Composing by the natural homomorphism ${\rm Cacl}(X)\to{\rm Cl}(X)$ from the Cartier divisor class group, we can define push-forward for all Cartier divisors.
@@ 3.3 Step 1: algebraic-geometric reductions
As $A$ is smooth, all Weil divisors on $A$ is Cartier. In particular, $f_{*}(R)$ is a Cartier divisor on $A$. It remains to show the ampleness. By [SER95, Proposition 1.4], it suffices to show that for any curve $C\subseteq A$, $f_{*}(R)\cdot C>0$. We first assume $R$ is ample and effective. After a suitable translation, we may assume that $f(R)$ and $C$ intersect properly. Then $C$ does not intersect any irreducible component of $f(R)$ of codimension greater than 1. If $f(R)\cap C=\emptyset$, for any curve $C^{\prime}\subseteq f^{-1}(C)$, we have $R\cap C^{\prime}=\emptyset$, which contradicts to the ampleness of $R$. Therefore $C\cdot f_{*}(R)>0$. This proves the case when $R$ is ample and effective.
For general $R$, there exists a positive integer $m$ such that $mR$ is linearly equivalent to $R_{1}+R_{2}$, where $R_{1}$ is ample and effective and $R_{2}$ is effective. By the argument above, $f_{*}(R_{1})$ is ample on $A$. Moreover, $f_{*}(R_{2})$ is effective by definition, so it is also nef by the group structure of $A$ (cf. [LAZ04, Example 1.4.7]). Therefore $f_{*}(R_{1})+f_{*}(R_{2})$ is ample on $A$. By [FUL98, Theorem 1.4], $f_{*}(R)$ is linearly equivalent to $f_{*}(R_{1})+f_{*}(R_{2})$, and thus it is also ample. ∎
@@
Since any effective divisor on an abelian variety is nef (cf. [LAZ04, Example 1.4.7]), we can easily prove that an effective divisor $D$ on an abelian variety is ample if and only if $D_{\rm red}$ is ample. Moreover, if $D_{1},D_{2}$ are two effective Cartier divisors on an abelian variety with ${\rm Supp}(D_{1})\subseteq{\rm Supp}(D_{2})$ and $D_{1}$ ample, then $D_{2}$ is also ample. This will be used in §3.4.
Basic setups
@@
Now we begin to prove Proposition 3.1 by contradiction. After taking normalization, we may assume that $X$ is normal. Assume to the contrary that the height of $\{x_{n}\}_{n\geq 1}$ is unbounded. After taking a subsequence, we can assume that the height of $\{x_{n}\}_{n\geq 1}$ tends to infinity as $n\to\infty$.
@@
First we repeat the following operations: Take $A_{1}=A,A_{0}={\rm Spec}\,K$ at the beginning. Then we have a finite map $f\colon X\to A_{1}\times A_{0}\times T_{K}$. In each step, if there is a nontrivial quotient abelian variety $A^{\prime}$ of $A_{1}$ (i.e. a surjective homomorphism $A_{1}\to A^{\prime}$) such that the height of the image of $\{x_{n}\}_{n\geq 1}$ in $A^{\prime}$ under the natural composition $X\to A_{1}\times A_{0}\times T_{K}\to A_{1}\to A^{\prime}$ does not tend to infinity, then after taking a subsequence of $\{x_{n}\}_{n\geq 1}$ we can assume that the image of $\{x_{n}\}_{n\geq 1}$ in $A^{\prime}$ has bounded height. Note that by our assumption in the previous paragraph, the height of $\{x_{n}\}_{n\geq 1}$ still tends to infinity. Choose an isogeny $A_{1}\overset{j}{\to}{\rm ker}(A_{1}\to A^{\prime})\times A^{\prime}$, compose $f$ with the map $j\times{\rm id}\colon A_{1}\times A_{0}\times T_{K}\to{\rm ker}(A_{1}\to A^{\prime})\times A^{\prime}\times A_{0}\times T_{K}$, and change $A_{1},A_{0}$ to ${\rm ker}(A_{1}\to A^{\prime}),A^{\prime}\times A_{0}$ respectively, we get a finite map $X\to A_{1}\times A_{0}\times T_{K}$. This ends a step. Since the dimension of $A_{1}$ decreases strictly, the operations terminate after finitely many steps. Then we get a finite map $X\to A_{1}\times A_{0}\times T_{K}$ such that the image of $\{x_{n}\}_{n\geq 1}$ in any nontrivial quotient abelian variety of $A_{1}$ has height tending to infinity, while the image of $\{x_{n}\}_{n\geq 1}$ in $A_{0}(K)$ has bounded height. Note that by the assumption in Proposition 3.1, the image of $\{x_{n}\}_{n\geq 1}$ in $T_{K}(K)$ is contained in $T(\mathbb{C})$, and thus also has bounded height. Combining these, we see that the abelian variety $A_{1}$ must be nontrivial.
@@
By Lang–Néron theorem (cf. [34, Theorem 3.4]), $V(A_{0},K)=(A_{0}(K)/A_{0}^{(K/\mathbb{C})}(\mathbb{C}))_{\mathbb{R}}$ is a finite-dimensional $\mathbb{R}$-vector space, and the canonical height on $A_{0}(K)$ (with respect to an ample line bundle on $A_{0}$) induces a positive definite quadratic form on $V(A_{0},K)$. As a consequence, there are only finitely many elements in $V(A_{0},K)$ with a fixed bounded height. Hence the image of $\{x_{n}\}_{n\geq 1}$ in $A_{0}(K)$ is contained in a finite union of translations of $A_{0}^{(K/\mathbb{C})}(\mathbb{C})$. As $\{x_{n}\}_{n\geq 1}$ is generic in $X$ and $X$ is irreducible, the image of $X$ in $A_{0}$ must be contained in a translation of $(A_{0}^{(K/\mathbb{C})})_{K}$. After composing $X\to A_{1}\times A_{0}\times T_{K}$ with a translation on $A_{0}$ (with identity on other components), we may assume that the image of $X$ in $A_{0}$ is contained in $(A_{0}^{(K/\mathbb{C})})_{K}$. Let $T^{\prime}$ be the Zariski closure of the image of $\cup_{n\geq 1}\{x_{n}\}$ in $A_{0}^{(K/\mathbb{C})}\times T$, which is a complex variety. Then the image of $X$ in $(A_{0}^{(K/\mathbb{C})})_{K}\times T_{K}$ is exactly $T^{\prime}_{K}$. In conclusion, we get a finite morphism $X\to A_{1}\times_{K}T^{\prime}_{K}$, whose composition with the projection $A_{1}\times_{K}T^{\prime}_{K}\to T^{\prime}_{K}$ is surjective.
@@
From now on we focus on the finite morphism $X\to A_{1}\times_{K}T^{\prime}_{K}$. For simplicity, we replace the original $A$ by $A_{1}$, replace the original $T$ by $T^{\prime}$, and replace the original $f$ by this map. In other words, we call this map $f\colon X\to A\times_{K}T_{K}$. After taking normalization for $T$, we may assume that $T$ is normal. Let $s_{n},t_{n}$ be the image of $x_{n}$ in $A(K),T_{K}(K)$ respectively, then $t_{n}\in T(\mathbb{C})$ is constant. Since $\{x_{n}\}_{n\geq 1}$ is generic in $X$ and $X\to T_{K}$ is surjective, $\{t_{n}\}_{n\geq 1}$ is also generic in $T$. Moreover, by our assumption, the image of $\{s_{n}\}_{n\geq 1}$ in any nontrivial quotient abelian variety of $A$ has height tending to infinity.
@@
We first consider the case when $f\colon X\to A\times_{K}T_{K}$ is not surjective. In this case, we can still use the arguments in [35] combining with a flattening process.